OpenStudy (kweb2017):

What is the equation (in slope-intercept form) of a line that passes through the point (3, -2) and has a slope of -2? y=2x+4 y=-2x+4 y+2=-2(x+3) 2x+y+4=0

pooja195 (pooja195):

Well y=mx+b is slope intercept form we know the slope y=-2x+b now take your coordinates and plug them in place of x and y -2=-2(3)+b solve for b :)

OpenStudy (kweb2017):

y=-2x+4?

pooja195 (pooja195):

Good :)

OpenStudy (kweb2017):

thank you <3

pooja195 (pooja195):

Welcome to openstudy :-) <3 and yw ^_^